# ESTIMATING RESPIRATORY EFFORT FROM NOCTURNAL BREATHING SOUNDS FOR OBSTRUCTIVE SLEEP APNOEA SCREENING

### v1

@@ Abstract
Obstructive sleep apnoea (OSA) is a prevalent condition with significant health consequences, yet many patients remain undiagnosed due to the complexity and cost of over-night polysomnography. Acoustic-based screening provides a scalable alternative, yet performance is limited by environmental noise and the lack of physiological context. Respiratory effort is a key signal used in clinical scoring of OSA events, but current approaches require additional contact sensors that reduce scalability and patient comfort. This paper presents the first study to estimate respiratory effort directly from nocturnal audio, enabling physiological context to be recovered from sound alone. We propose a latent-space fusion framework that integrates the estimated effort embeddings with acoustic features for OSA detection. Using a dataset of 157 nights from 103 participants recorded in home environments, our respiratory effort estimator achieves a concordance correlation coefficient of 0.48, capturing meaningful respiratory dynamics. Fusing effort and audio improves sensitivity and AUC over audio-only baselines, especially at low apnoea–hypopnoea index thresholds. The proposed approach requires only smartphone audio at test time, which enables sensor-free, scalable, and longitudinal OSA monitoring.
@@ 1 Introduction
The diagnostic reference test for OSA is overnight polysomnography (PSG), which typically employs more than 12 sensors. Sleep clinicians manually score OSA events from these multimodal signals and compute the apnoea-hypopnoea index (AHI), defined as the number of events per hour of sleep. Although accurate, PSG is costly, labour-intensive, and restricted to specialised laboratories. The use of multiple sensors can disrupt sleep, and long waiting lists further delay diagnosis. Consequently, a large proportion of adults with OSA remain undiagnosed or untreated [5].
Recent evidence further highlights the limitations of one-night PSG assessments, with AHI demonstrating high night-to-night variability [6, 7]. This is clinically significant given its associations with other conditions. Longitudinal OSA monitoring provides a more reliable picture of habitual disease burden, yet current clinical pathways rarely enable frequent measurement.
Acoustic-based monitoring has been shown to be a promising non-invasive alternative. Snoring and breathing sounds carry physiological information about airway obstruction, and several studies have demonstrated feasibility of acoustic-based OSA screening. For example, Nakano et al. [8] reported strong performance for detecting mild/moderate/severe cases (AUC 0.99), using tracheal microphones in controlled hospital environments. In contrast, [9] investigated smartphone-based acoustic monitoring at home, which achieved competitive though lower AUCs of about 0.7 for mild and 0.8 for moderate/severe OSA. Similar declines in performance in home environments were reported in [10], who conducted a noise simulation and found acoustic-based OSA events detection is sensitive to specific noises including air conditioning and the human voice. To improve the robustness of audio-based methods, many studies introduce other physiological signals or demographic data as additional information [11, 12, 13]. Respiratory effort, in particular, has been shown to provide complementary information to acoustic features and significantly improve OSA screening [12]. However, current methods require additional sensors to capture effort, which limits scalability, increases cost, and reduces patient acceptability in home or longitudinal monitoring scenarios.
@@
The rest of the paper is organised as follows. Section 2 describes the dataset used in this study. Section 3 presents the proposed system for estimating respiratory effort from acoustics and integrating these embeddings for OSA detection. Section 4 describes the experimental setup, and results are reported and discussed in Section 5. Finally, Section 6 concludes the paper and outlines future directions.
@@ 3.1 Acoustic-based respiratory effort estimation
The proposed respiratory effort estimator consists of three main components: a convolutional neural network (CNN) feature extractor, an Long short-term memory (LSTM) encoder, and a decoder with interpolation. The CNN feature extractor uses smaller pooling kernels than the audio-only baseline to preserve temporal resolution, converting log-Mel spectrograms from (1500 $\times$ 64) to (187 $\times$ 128) (time frames $\times$ embedding dimensions). The LSTM encoder then processes these 187 frames to produce hidden vectors that capture temporal respiratory dynamics. For respiratory effort prediction, a linear decoder projects each hidden vector to a single value, generating a 187-point sequence.
@@ 3.2 SDB events detection and AHI estimation
Specifically, the respiratory embeddings are obtained by averaging the LSTM hidden states across time, producing a fixed-dimensional representation. This is concatenated with the audio embeddings from a baseline audio-only encoder, and the combined representation is used to predict the occurrence of apnoea and hypopnoea events. Finally, predicted events occurring in close succession are merged to form OSA episodes, from which night-level AHI is derived.
The AHI value is calculated as:
@@ 4 EXPERIMENTAL SETUP
4.1 Models for comparison
We first implemented an audio-only baseline using a similar CNN backbone and training procedure but without respiratory inputs, following [12]. The model comprised three convolutional blocks with 16, 32, and 64 filters, each followed by batch normalisation, ReLU activation, and max-pooling. The resulting feature maps were flattened, projected to a 512-dimensional vector, and then passed through a final fully connected layer with sigmoid activation to produce single logit.
We then evaluated our proposed latent-space fusion (LSF) model, which concatenated a frozen respiratory embedding (inferred from audio) with the audio-encoder output and passed the fused vector through a fusion layer and an MLP for classification.
For context, we also report an oracle results from prior work on the same dataset [12] that uses reference respiratory effort, representing an upper bound on performance.
@@ 4.3 Evaluation metrics
The performance of respiratory effort signal estimator was evaluated at 30-s segment level with CCC, MAE and RMSE. Mean and standard deviation across all 10 folds are reported for each metric.
We used AHI cut-offs of 5, 15, and 30 events/h to stratify OSA severity groups (healthy, mild, moderate, severe). These thresholds are widely adopted for diagnosis and treatment of OSA [16]. An AHI $\geq$ 5 indicates OSA, and AHI $\geq$ 15 tends to lead to more arousals [17]. Untreated severe OSA (over 30) can lead to conditions such as hypertension and heart failure [18]. For each cut-offs, we evaluate the sensitivity, specificity and the area under the curve (AUC).
@@ 5.1 Respiratory effort prediction from audio
Table 2 reports the performance of respiratory effort prediction from nocturnal audio. Averaged across all folds, the proposed estimator achieved a CCC of 0.478, RMSE of 1.053, and MAE of 0.793. This suggest that audio carries information about underlying respiratory effort dynamics.
@@
Fig. 2 illustrates a typical example of predicted respiratory effort from audio compared to the ground truth measured effort signal. In the first example, the estimated effort capture the overall trend of OSA events shown in the effort, with clear flat regions during the pause of snore sounds. Although the CCC correlation value is only around 0.48, the predicted effort successfully captures the physiological effort dynamics. The second example shows a close alignment with ground truth, while the third example highlights the situation of phase misalignment, where effort trends are reproduced but shifted in time. For segments lacking audible breathing, the predicted effort signals are nearly a flat line with little fluctuation. Overall, these results suggest that despite noise and variability, audio can provide a meaningful proxy for respiratory effort.
Related work using clean speech recordings on a smaller dataset (25 hours, 26 speakers) reported higher accuracy, with CCC of 0.6 and RMSE of 0.15 [19]. In contrast, our results reflect the challenges of a larger, noisier sleep sound dataset collected from real-world home environments using distant smartphone microphones. Higher MAE and RMSE may also stem from downsampling through pooling layers, which reduces temporal resolution of the effort signal. Additionally, temporal misalignment between audio and respiratory signals can lead to occasional low correlations, as illustrated in the bottom plot in Fig. 2.
@@ 5.2 OSA severity classification
AHI 5 15 30
@@
Latent Fusion – measured effort [12] Sensitivity 0.81 0.81 0.88
@@
Latent Fusion – estimated effort embeddings Sensitivity 0.88 0.83 0.78
@@
Table 3: OSA severity classification results at different AHI cut-off thresholds across models.
Table 3 lists the OSA severity classification results at different AHI cut-off thresholds across models. At the lowest AHI cut-off (5 events/h), the proposed LSF model with estimated respiratory effort embeddings achieves the highest sensitivity (0.88), slightly outperforming both the audio-only baseline (0.86) and the oracle system with measured effort (0.81). This suggests that incorporating estimated effort provides additional physiological context that is especially valuable for detecting mild OSA, where events are less pronounced and more difficult to identify from audio alone. However, this improvement in sensitivity comes at the expense of specificity, which drops compared to the oracle system (0.69 vs. 0.82), but still better than the audio-only baseline (0.62).
At the moderate AHI cut-offs (15 events/h), LSF with estimated effort again improves sensitivity over the baseline (0.83 vs. 0.81) while maintaining competitive specificity (0.78). The resulting AUC (0.88) is the highest across all systems, indicating that estimated effort embeddings contribute most strongly at clinically relevant mid-range severities where distinguishing between moderate and severe OSA is critical.
@@
These results suggest that audio alone contains latent physiological cues that can support effective at-home OSA screening. The modest overall gains can be explained by several factors: (i) limited data and strong variability in smartphone recordings and real-world room acoustics, (ii) potential misalignment between audio and effort during zero-lag evaluation, and (iii) the coarseness of categorical AHI cut-offs compared to the continuous respiratory representation.
Overall, these results demonstrate that the proposed system, despite relying on noisy estimated effort signals inferred from audio, achieves AUC values close to those of models using measured effort. The key advantage of the proposed method is its scalability and non-invasiveness: at test time, only smartphone audio is required, which avoids additional sensors that may reduce patient acceptability or limit longitudinal monitoring.
@@ 6 Conclusion
In this study, we proposed a latent-space fusion framework that exploits respiratory effort embeddings inferred directly from nocturnal audio to improve OSA screening. Our experiments demonstrate that audio-inferred effort captures meaningful respiratory dynamics and, when fused with acoustic features, provides modest but consistent gains in sensitivity and AUC at clinically relevant AHI thresholds. Compared with prior work using measured respiratory effort [12], our approach achieves smaller improvements, reflecting the challenges of noise, temporal misalignment, and reduced precision in predicted effort signals.
Future work will explore simpler physiological targets, such as breathing-rate variability, to guide the audio-only classifier [20]. We also plan to improve effort prediction through domain-adapted snore/breathing foundation models, robust noise- and device-augmentation, and lag-tolerant training objectives. To employ respiratory acoustic foundation models [21, 22], additional fine-tuning to reduce the domain shift would be needed.

### v2

@@ Abstract
Obstructive sleep apnoea (OSA) is a prevalent condition with significant health consequences, yet many patients remain undiagnosed due to the complexity and cost of overnight polysomnography. Acoustic-based screening provides a scalable alternative, yet performance is limited by environmental noise and the lack of physiological context. Respiratory effort is a key signal used in clinical scoring of OSA events, but current approaches require additional contact sensors that reduce scalability and patient comfort. This paper presents the first study to estimate respiratory effort directly from nocturnal audio, enabling physiological context to be recovered from sound alone. We propose a latent-space fusion framework that integrates the estimated effort embeddings with acoustic features for OSA detection. Using a dataset of 157 nights from 103 participants recorded in home environments, our respiratory effort estimator achieves a concordance correlation coefficient of 0.48, capturing meaningful respiratory dynamics. Fusing effort and audio improves sensitivity and AUC over audio-only baselines, especially at low apnoea–hypopnoea index thresholds. The proposed approach requires only smartphone audio at test time, which enables sensor-free, scalable, and longitudinal OSA monitoring.
@@ 1 Introduction
The diagnostic reference test for OSA is overnight polysomnography (PSG), which typically employs more than 12 sensors. Sleep clinicians manually score OSA events from these multimodal signals to compute the apnoea-hypopnoea index (AHI). This metric aggregates both aponeas (complete airway collapses) and hypopnoeas (partial airway collapses) equally, defining the AHI as the total number of combined events per hour of sleep.   Although accurate, PSG is costly, labour-intensive, and restricted to specialised laboratories. The use of multiple sensors can disrupt sleep, and long waiting lists further delay diagnosis. Consequently, a large proportion of adults with OSA remain undiagnosed or untreated [17]. Recent evidence further highlights the limitations of one-night PSG assessments, with AHI demonstrating high night-to-night variability [12, 10]. This is clinically significant given its associations with other conditions. Longitudinal OSA monitoring provides a more reliable picture of habitual disease burden, yet current clinical pathways rarely enable frequent measurement.
Acoustic-based monitoring has been shown to be a promising non-invasive alternative. Snoring and breathing sounds carry physiological information about airway obstruction, and several studies have demonstrated feasibility of acoustic-based OSA screening. These studies typically address one of three tasks: SDB event detection (locating apnoea and hypopnoea events), AHI estimation (calculating the index from detected events), or severity classification (categorising subjects into healthy/mild/moderate/severe groups based on AHI cut-offs presented in Table 1). For severity classification, Nakano et al. [14] reported strong performance for detecting mild/moderate/severe cases (AUC 0.99), using tracheal microphones in controlled hospital environments. In contrast, [15] investigated smartphone-based acoustic monitoring at home, which achieved competitive though lower AUCs of about 0.7 for mild and 0.8 for moderate/severe OSA. Similar declines in performance in home environments were reported in [9], who conducted a noise simulation and found that acoustic-based OSA events detection is sensitive to specific noises including air conditioning and the human voice. To improve the robustness of audio-based methods, many studies introduce other physiological signals or demographic data as additional information [1, 16, 11]. Respiratory effort, in particular, has been shown to provide complementary information to acoustic features and significantly improve OSA screening [16]. However, current methods require additional sensors to capture effort, which limits scalability, increases cost, and reduces patient acceptability in home or longitudinal monitoring scenarios.
@@ 3.1 Acoustic-based respiratory effort estimation
The proposed respiratory effort estimator consists of three main components: a convolutional neural network (CNN) feature extractor, a long short-term memory (LSTM) encoder, and a decoder with interpolation. The CNN feature extractor uses smaller pooling kernels than the audio-only baseline to preserve temporal resolution, converting log-Mel spectrograms from 1500 $\times$ 64 to 187 $\times$ 128 (time frames $\times$ embedding dimensions). The LSTM encoder then processes these 187 frames to produce hidden vectors that capture temporal respiratory dynamics. For respiratory effort prediction, a linear decoder projects each hidden vector to a single value, generating a 187-point sequence.
@@ 3.2 SDB events detection and AHI estimation
Specifically, the respiratory embeddings are obtained by averaging the LSTM hidden states across time, producing a fixed-dimensional representation. This is concatenated with the audio embeddings from a baseline audio-only encoder, and the combined representation is used to predict the occurrence of apnoea and hypopnoea events. Finally, predicted events occurring in close succession are merged to form OSA episodes, from which night-level AHI is derived. The AHI value is calculated as:
@@ 4 EXPERIMENTAL SETUP
4.1 Baselines
We first implemented an audio-only baseline that mirrors the backbone architecture and training procedure of our proposed model, but without any respiratory-related inputs, following [16]. The network consists of three convolutional blocks with 16, 32, and 64 filters, respectively. Each block comprises convolution, batch normalisation, ReLU activation, and max-pooling. The resulting feature maps are flattened and projected to a 512-dimensional embedding, which is passed through a final fully connected layer with sigmoid activation to produce a single output score.
For reference, we also report results from an oracle system that directly uses ground-truth respiratory effort signals, as reported in [16] on the same dataset. This oracle model represents an upper bound on performance achievable when high-quality respiratory measurements are available at inference time.
Finally, we evaluated our proposed latent-space fusion (LSF) model. In this setting, a frozen respiratory embedding inferred from audio is concatenated with the output of the audio encoder. The fused representation is then passed through a fusion layer followed by a multi-layer perceptron (MLP) to produce the final classification output.
@@ 4.3 Evaluation metrics
The performance of the respiratory effort signal estimator was evaluated at 30-s segment level with CCC, MAE and RMSE. Mean and standard deviation across all 10 folds are reported for each metric. We used AHI cut-offs of 5, 15, and 30 events/h to stratify OSA severity groups (healthy, mild, moderate, severe). These thresholds are widely adopted for diagnosis and treatment of OSA [4]. An AHI $\geq$ 5 indicates mild OSA, and AHI $\geq$ 15 tends to lead to more arousals [2]. Untreated severe OSA (over 30) can lead to conditions such as hypertension and heart failure [3]. For each cut-off, we evaluate the sensitivity, specificity and the area under the curve (AUC).
@@ 5.1 Respiratory effort prediction from audio
Table 2 summarises the performance of respiratory effort prediction from nocturnal audio. Averaged across all cross-validation folds, the proposed estimator achieved a CCC of 0.478, with an RMSE of 1.053 and an MAE of 0.793. These results indicate that nocturnal audio contains informative cues related to underlying respiratory effort dynamics, despite the absence of direct physiological measurements.
@@
Fig. 2 presents representative examples of predicted respiratory effort signals compared with the corresponding ground-truth measurements. In the first example, the estimated effort captures the overall temporal trend of respiratory events, including prolonged low-activity regions corresponding to pauses in snore-related sounds. Although the overall CCC is moderate, the predicted signal reproduces key physiological patterns. The second example shows close temporal alignment with the ground truth, whereas the third illustrates a case of phase misalignment, where effort dynamics are preserved but shifted in time. In segments with minimal audible breathing activity, the predicted effort remains largely flat, reflecting the lack of informative acoustic cues. Overall, these examples demonstrate that audio-derived estimates can serve as a meaningful, albeit imperfect, proxy for respiratory effort.
Prior work using clean speech recordings on a smaller, controlled dataset (25 hours, 26 speakers) reported higher prediction accuracy, with a CCC of 0.6 and an RMSE of 0.15 [13]. In contrast, the present results reflect the increased difficulty of modelling respiratory effort from a substantially larger and noisier sleep sound dataset collected in real-world home environments using distant smartphone microphones. The relatively higher MAE and RMSE may also be partially attributed to temporal downsampling introduced by pooling operations in the network, which reduces the temporal resolution of the predicted effort signal. Furthermore, residual temporal misalignment between audio and respiratory measurements can lead to reduced correlation in some segments, as illustrated in the lower example of Fig. 2.
@@ 5.2 OSA severity classification
Table 3 summarises OSA severity classification performance at three clinically relevant AHI cut-offs (5, 15, and 30 events/h). Results are aggregated across all cross-validation folds to reflect performance on the full dataset, as the limited number of nights per fold makes fold-wise variance estimates unstable and difficult to interpret.
At the lowest AHI threshold (5 events/h), corresponding to detection of mild OSA, the proposed LSF model with estimated respiratory effort embeddings achieves the highest sensitivity (0.88), outperforming both the audio-only baseline (0.86) and the oracle system using measured effort (0.81). This indicates that incorporating estimated effort provides complementary physiological information that is particularly valuable when OSA events are subtle and less acoustically salient. This gain in sensitivity is accompanied by a reduction in specificity compared to the oracle model (0.69 vs. 0.82), though specificity remains higher than that of the audio-only baseline (0.62).
AHI Cut-off 5 15 30
@@
Latent Fusion with Reference Effort [16] Sensitivity 0.81 0.81 0.88
@@
Proposed LSF Model Sensitivity 0.88 0.83 0.78
@@
Table 3: OSA severity classification results at various clinical AHI cut-offs (5, 15, 30 events/h). Results are aggregated across all test folds to represent the full dataset. The second row shows the class distribution between Nights << Cut-off | Nights $\geq$ Cut-off.
At the moderate AHI threshold (15 events/h), the proposed LSF model again improves sensitivity over the audio-only baseline (0.83 vs. 0.81) while maintaining competitive specificity (0.78). In contrast, the audio-only baseline exhibits less stable performance, with sensitivity varying substantially across folds and dropping to as low as 0.72 in some splits. This behaviour suggests that the auxiliary effort estimation task acts as a form of regularisation, encouraging the model to learn physiologically meaningful representations rather than overfitting to dataset-specific acoustic patterns. The proposed model achieves an AUC of 0.88, the highest among all models at this mid-range threshold, where accurate discrimination between moderate and severe OSA is essential for treatment decisions.
@@
Overall, these results demonstrate that nocturnal audio alone contains latent physiological information sufficient to support effective at-home OSA screening. The modest but consistent gains achieved by the proposed LSF model can be attributed to several factors: (i) limited data and substantial variability in smartphone recordings and real-world room acoustics, (ii) residual temporal misalignment between audio and effort signals during zero-lag evaluation, and (iii) the use of categorical AHI thresholds, which may not fully capture the benefits of continuous respiratory representations.
Despite these challenges, the proposed method achieves AUC values close to those of systems relying on directly measured reference respiratory effort. Its key advantage lies in scalability and practicality: at inference time, only smartphone audio is required, which avoids additional sensors that may reduce patient acceptability or hinder long-term, large-scale monitoring.
@@ 6 Conclusion
We presented a latent-space fusion framework that integrates respiratory effort embeddings inferred from nocturnal audio to enhance audio-only OSA screening. Our experiments demonstrate that audio-inferred effort captures meaningful respiratory dynamics and, when fused with acoustic features, provides modest but consistent gains in sensitivity and AUC at clinically relevant AHI thresholds. Compared with prior work using measured respiratory effort [16], our approach achieves smaller improvements, reflecting the challenges of noise, temporal misalignment, and reduced precision in predicted effort signals.
A key limitation is the lack of a parameter-matched ablation analysis, which prevents a separation of the benefits of effort representations from model capacity. Preliminary experiments indicated that removing effort-estimation pre-training led to substantial overfitting in the proposed LSF framework at low AHI cut-offs. This suggests that the performance improvements were due to the benefits of learning an auxiliary task not increased model capacity. However, this remains to be confirmed in future studies. Future work will also report additional evaluation metrics, including precision and F1-score, to better characterise the clinical utility of our approach.